OpenStudy (anonymous):

Determine algebraically whether the given function is even odd or neither. f(x) = 3x + l4xl.

OpenStudy (anonymous):

Welcome to OpenStudy! Message me if you have any questions!

OpenStudy (anonymous):

plug in -x for x and if you get the same function then it is even, if it's the same function with a factor of -1, then it's odd, if neither then neither.

OpenStudy (anonymous):

def: \[f(x) = x\] is even, and, \[f(-x) = -x\] is odd. using this definition we plug into the original function and test \[f(1) = -3 + 4 = 1\] and \[f(-1) = 3 + 4 = 7\] ...based on southpaw's statement, what can you conclude.